A collector is using a magnifying lens with a radius of curvature of 21.0 cm to examine the finer details of a butterfly with a magnification of 5.10. At what distance between the butterfly and the lens does she achieve the desired magnification?

Group of answer choices

13.1 cm

12.6 cm

8.40 cm

8.80 cm

16.8 cm
Answer #1

> M=/ tensmaler formula * = (414,- 62) at & RU RE-R s.id u=-d; V=59d; f=.421ems J= R = 22cm traz lens formula -1 +5.1 = 1/2